OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS Roger is 13 years younger than his sister Jackie. Together the sum of their ages of 91. How old is both Roger and Jackie?

OpenStudy (lgbasallote):

let x = jackie's age x-13 = roger's age x + (x-13) = 91 2x - 13 = 91 2x = 106 x = 53 x - 13 = 40 so Jackie is 53 years old and Roger is 40 years old.

OpenStudy (lgbasallote):

sorry that was 2x = 104 x = 52 x-13 = 39 Jackie = 52 yrs old Roger = 39 yrs old

OpenStudy (anonymous):

so in this equation there is no y=?

OpenStudy (lgbasallote):

well y is just a variable. You can also use y. For example y = Jackie's age y-13 = Roger's age then you repeat the whole process.

OpenStudy (anonymous):

what i ment was you dont need both y and x in the equation? y = roger x = jackie?

OpenStudy (lgbasallote):

oh. i see your point. okay, i'll do it your way. x = Jackie y = Roger y = x - 13 x+ y = 91 using systems of linear eqautions, -x + y = -13 x+y = 91 add the two equations... 2y = 78 y = 39 39 = x - 13 x = 29 + 13 x = 52 was that what you meant?

OpenStudy (anonymous):

yes thankyou so much!